# Massive deformations of supersymmetric Yang-Mills matrix models

### v1

@@ 3 Classification of massive SYM matrix models
In this section, we provide the main idea that allows to classify massive SYM matrix models. Here we will assume that we have one Majorana or Majorana-Weyl spinor, meaning $D=3,4,10$. In $D=6$, we have more than one Majorana spinor so we will treat this case separately (see appendix A.2). [^11] Note that the $D=2$ pure SYM matrix model does not allow any mass-deformation. This is because one can neither write a fermionic mass term nor a Myers term in $D=2$. The constraints (24) and (25) then imply that all deformed quantities have to vanish.
Consider a massive deformation of $S_{0}$ (3) parameterized by a mass-parameter $\mu$,
@@
while there exist no (polynomial) combination that can give lower dimensional operators. [^12] Note that $\mathrm{Tr}X^{I}$ vanishes for traceless matrices. On the other hand, if one allows the matrices to have a trace, the term $\mathrm{Tr}X_{I}$ can always be removed by completing the square with the trace contribution coming from $S_{IJ}\mathrm{Tr}X_{I}X_{J}$. Inspired by Myers:1999ps , we refer to $S_{IJK}\mathrm{Tr}X^{I}X^{J}X^{K}$ as the Myers term.
The supersymmetries (4) may also need to be deformed in order to leave (15) invariant for all $\epsilon$. We allow the generic form
@@
Note that $\epsilon$ has mass dimension $-1/2$. Thus, $\delta_{B}^{(1)}X^{I},\delta_{B}^{(2)}X^{I},\ldots$ must be fermionic operators of dimension $1/2,-1/2,\ldots$. Since our fermions have dimension $3/2$, we cannot build such operators and thus $\delta_{B}^{(k)}=0$. On the other hand, $\delta_{F}^{(1)}\psi$ must be a bosonic operator of dimension 1. The only possibility is $\delta_{F}^{(1)}\psi\sim X_{I}$ while $\delta_{F}^{(k>1)}=0$. Putting everything together, the most general possibility at this point is
$\begin{split}S=S_{0}+\frac{1}{g_{\mathrm{YM}}^{2}}\mathrm{Tr}\Biggl{\{}\mu\bar{\psi}M\psi+\mu S_{IJK}X^{I}X^{J}X^{K}+\mu^{2}S_{IJ}X^{I}X^{J}\Biggr{\}}\,,\end{split}$ (19)
@@
$\begin{split}\delta_{\epsilon}X^{I}&=\bar{\epsilon}\Gamma^{I}\psi\,,\\ \delta_{\epsilon}\psi&=\frac{i}{2}\Gamma^{IJ}\epsilon[X_{I},X_{J}]+\mu F^{I}\epsilon X_{I}\,,\end{split}$ (20)
@@
$\begin{split}0=\mathrm{Tr}\Biggl{\{}[\bar{\psi},X^{I}](i\Gamma^{I}F^{J}\epsilon-iM\Gamma^{IJ}\epsilon)X^{J}+3S_{KIJ}\bar{\psi}X^{I}\Gamma^{K}\epsilon X^{J}\Biggr{\}}\,,\end{split}$ (22)
@@
This needs to be imposed for all spinors $\epsilon$, which are Majorana or Majorana-Weyl depending on the dimension. In order to proceed solving these equations, one needs to specify the dimension $D$. The detailed analysis is carried out for each $D$ in appendix A. The resulting models are discussed in the next sections and summarized in table 1.
@@ 4 Ten-dimensional matrix model: polarized IKKT
The polarized IKKT model, originally introduced in Bonelli_2002 and recently examined in Hartnoll:2024csr ; Komatsu:2024bop ; Komatsu:2024ydh ; Hartnoll:2025ecj , is a mass-deformed version of the IKKT model Ishibashi_1997 that retains invariance under sixteen (deformed) supersymmetries. We consider this model as the most interesting among all the matrix models, because of its richness as a holographic description of ten-dimensional geometries Komatsu:2024bop ; Komatsu:2024ydh ; Hartnoll:2024csr . In appendix A.1, we show that it is the unique [^13] Up to redefinition of $\mu$ and $SO(10)/(SO(3)\times SO(7))$ rotations. Moreover, we restrict to real bosonic deformations. mass-deformation of IKKT that has sixteen supersymmetries and preserves $SU(N)$ gauge invariance.
@@
$\begin{split}\delta_{\epsilon}X^{I}&=\bar{\epsilon}\Gamma^{I}\psi\,,\\ \delta_{\epsilon}\psi&=\frac{i}{2}\Gamma^{IJ}\epsilon[X_{I},X_{J}]+\frac{3\mu}{8}\Gamma^{123}\Gamma^{i}\epsilon X_{i}+\frac{\mu}{8}\Gamma^{123}\Gamma^{p}\epsilon X_{p}\,,\end{split}$ (27)
where $\epsilon$ is a Majorana-Weyl spinor (commuting) parameter, and $\delta_{\epsilon}$ is an anti-commuting (Grassmann-odd) operator. The supersymmetries together with the global $SO(3)\times SO(7)$ form either a gauged $F(4)$ or a gauged $F^{1}(4)$ superalgebra in Lorentzian signature. [^14] The anticommutator of two supersymmetries takes exactly the form expected from $F(4)$ as was shown in Komatsu:2024ydh . Note that as in IKKT, $g_{\mathrm{YM}}$ can be scaled out of the integral, by rescaling $X_{I}=\sqrt{g_{\mathrm{YM}}}\tilde{X}_{I}$ and $\psi_{\alpha}=g_{\mathrm{YM}}^{3/4}\tilde{\psi}_{\alpha}$. This leaves only $\Omega\equiv\mu/\sqrt{g_{\mathrm{YM}}}$ as a mass-parameter.
@@
Let us now discuss observables, in dimensionless units where the only tunable parameter is $\Omega\equiv\mu/\sqrt{g_{\mathrm{YM}}}$. Firstly, the fixed $N$, $\Omega\to 0$ limit is accessible analytically (at leading order). In this case, one can integrate out the off-diagonal components. This allows to compute a large set of gauge invariant observables in the strict massless limit as shown in (220). In particular, this analysis yields
@@
The large $\Omega$ behaviour $\Omega\gg N^{1/4}$ was obtained in Hartnoll:2024csr yielding
@@ 5 Six-dimensional matrix models
The supersymmetry transformations take the form [^15] Recall the we use the convention where $\epsilon$ is a commuting parameter while $\delta_{\epsilon}$ is an anticommuting operator.
@@
In appendix A.2 we derive all the mass-deformations of $S_{0}^{(D=6)}$ that carry eight supersymmetries as well as $SU(N)$ gauge invariance. We find two different models that we will denote by type I and type II. The type I model breaks $SO(6)\to SO(3)\times SO(3)$, and possesses a Myers term, so that the matrix integral has many different saddles as in polarized IKKT. The type II model has $SO(5)$ global symmetry and only has a trivial saddle.
@@ 5.1 Type I: $SO(6)\to SO(3)\times SO(3)$
Note that the $SU(2)$ symmetry $\psi_{a}\to U_{ab}\psi_{b}$ is preserved. In Lorentzian signature, the total global symmetry group would thus be $SO(2,1)\times SO(3)\times SU(2)$, and we have 8 supercharges. We thus expect the superalgebra to lie in the exceptional family $D(2,1,\alpha)$ VanProeyen:1999ni . This is indeed the case as we show in appendix A.2.1, where we find the specific value $\alpha=1/2$ which is not isomorphic to $OSp(4|2)$. [^16] It is surprising that we did not find models with arbitrary values of $\alpha$, since $\alpha=1/2$ seems arbitrary. However, it is not hard to see by allowing arbitrary values of the coefficients in (37) and (38) that supersymmetry is an algebraic constraint that leaves no free parameter except $\mu$. Thus, $\alpha=1/2$ is, surprisingly, the unique possibility.
@@
Finally, let us discuss observables in dimensionless coordinates $(g_{\mathrm{YM}}\to 1,\mu\to\Omega\equiv\mu/\sqrt{g_{\mathrm{YM}}})$. In the strict massless limit at fixed $N$, we can integrate out off-diagonals as shown in appendix C. This allows to compute a large class of observables, including
@@ 5.2 Type II: $SO(6)\to SO(5)$
$\begin{split}S_{D=6,\mathrm{type\ II}}=\frac{1}{g_{\mathrm{YM}}^{2}}\operatorname{Tr}\Biggl{[}-\frac{1}{4}&[X_{I},X_{J}]^{2}-\frac{i}{2}\epsilon^{ab}\bar{\psi}_{b}\Gamma^{I}[X_{I},\psi_{a}]\\ &+\mu^{2}X_{i}X_{i}+3\mu^{2}X_{6}^{2}+2\mu\bar{\psi}_{2}\Gamma^{12345}\psi_{1}\Biggr{]}\,,\end{split}$ (45)
@@
Note that the $SU(2)$ symmetry $\psi_{a}\to U_{ab}\psi_{b}$ with $U\in SU(2)$ remains a symmetry only if $U^{\top}sU=s$ where $s$ is a two by two symmetric matrix with components $s_{12}=s_{21}=1$, $s_{11}=s_{22}=0$. This restricts it to a one parameter symmetry, infinitesimally equivalent to a $U(1)$. In Lorentzian signature, we thus have either $SO(5)\times U(1)$ or $SO(4,1)\times U(1)$ as the symmetry group, together with 8 supercharges. As we show in appendix A.2.2, the corresponding superalgebra is either gauged $OSp^{*}(2|4)$ or gauged $OSp^{*}(2|2,2)$ respectively.
Let us now discuss observables in dimensionless units $(g_{\mathrm{YM}}\to 1,\mu\to\Omega\equiv\mu/\sqrt{g_{\mathrm{YM}}})$. In the strict massless limit at fixed $N$, we can integrate out off-diagonals as shown in appendix C. This allows to compute a large class of observables, including
@@ 6 Four-dimensional matrix models
where $\mathbf{X}_{I}$ are $(N^{2}-1)\times(N^{2}-1)$ matrices written in the adjoint representation. Namely, the original matrices are expanded as $X_{I}=X_{I}^{A}T^{A}$ where $T^{A}$ are the $SU(N)$ generators in the fundamental representation normalized by $\mathrm{Tr}T^{A}T^{B}=\delta^{AB}$. The structure constants are defined by $[T_{A},T_{B}]=if_{ABC}T_{C}$. Finally, we define the $(N^{2}-1)\times(N^{2}-1)$ matrices $\mathbf{X}_{I}$, by the components $(\mathbf{X}_{I})_{AB}\equiv f_{ABC}X_{I}^{C}$. Thus, they are real and antisymmetric. It is important to note that contrary to all other dimensions, this determinant is always positive semi-definite. The proof was first established in Ambjorn:2000bf . We review it in appendix B.
In appendix A.3, we derive all the mass-deformations of the $D=4$ pure SYM matrix model that carry four supersymmetries as well as $SU(N)$ gauge invariance. We find two models that we denote by type I and type II. The type I model preserves $SO(4)$ symmetry, and is parameterized by a single mass-parameter $\mu$. The type II model breaks $SO(4)\to SO(3)$, and it is parameterized by a continuous family of two mass-parameters $\mu_{1},\mu_{2}$. It has a Myers term, thus the matrix integral has many saddles as in polarized IKKT. The Pfaffians get deformed to functions of $\mu$ (type I) or $\mu_{1}$ (type II). As long as $\mu,\mu_{1}\in\mathbb{R}$, the non-negativity of the Pfaffian is preserved.
@@ 6.1 Type I: $SO(4)\to SO(4)$
In Lorentzian signature, the symmetry group would be $SO(3,1)$, and there are 4 supercharges. In appendix A.3.1, we show that the resulting superalgebra is a (gauged) (complexified) direct sum of two simple Lie superalgebras $OSp(1|2)\oplus OSp(1|2)$. [^18] This is the same spirit as $\mathfrak{so}(3,1)\cong\mathfrak{sl}(2,\mathbb{R})\oplus_{\mathbb{C}}\mathfrak{sl}(2,\mathbb{R})$.
@@
The normalization for the measure is explained in appendix E. The matrices $\mathbf{X}_{I}$ were defined below (55). As we show in appendix B, the Pfaffian is always positive semi-definite when $\Omega\in\mathbb{R}$. This implies that it can be written as the square root of a determinant. Moreover, this determinant can be simplied, yielding
@@
$\begin{split}Z_{SU(N)}\xrightarrow{\Omega\to 0}\frac{(2\pi)^{2(N^{2}-1)}}{\prod_{k=1}^{N}(k!)}\left(\frac{1}{\Omega^{2}}\right)^{N-1}\,,\\ \langle\mathrm{Tr}X_{1}^{2p}\rangle_{SU(N)}\xrightarrow{\Omega\to 0}N\left(\frac{N-1}{N\Omega^{2}}\right)^{p}\frac{\Gamma(p+\frac{1}{2})}{\sqrt{\pi}}\,,\end{split}$ (62)
@@ 6.2 Type II: $SO(4)\to SO(3)$
$\begin{split}S_{D=4,\mathrm{type\ II}}=\frac{1}{g_{\mathrm{YM}}^{2}}\operatorname{Tr}&\Biggl{[}-\frac{1}{4}[X_{I},X_{J}]^{2}-\frac{i}{2}\bar{\psi}\Gamma^{I}[X_{I},\psi]+i\frac{2}{3}(\mu_{1}+\mu_{2})\epsilon_{ijk}X_{i}X_{j}X_{k}\\ &+\left(\frac{2}{9}\mu_{1}^{2}+\frac{1}{3}\mu_{1}\mu_{2}\right)X_{i}X_{i}+\frac{1}{3}\mu_{1}\mu_{2}X_{4}^{2}-\frac{1}{3}\mu_{1}\bar{\psi}\Gamma^{123}\psi\Biggr{]}\,,\end{split}$ (64)
@@
In addition to these symmetries, there are chiral rotations $\psi\to e^{i\alpha\Gamma_{*}}\psi$ that form a $U(1)$. In Lorentzian signature, the symmetry group is either $SO(3)\times U(1)$ or $SO(2,1)\times U(1)$, and there are 4 supercharges. In appendix A.3.2, we show that the corresponding superalgebra is either gauged $SU(2|1)$ or gauged $SU(1,1|1)$ respectively VanProeyen:1999ni .
@@
In dimensionless units $(g_{\mathrm{YM}}\to 1,\mu_{i}\to\Omega_{i}\equiv\mu_{i}/\sqrt{g_{\mathrm{YM}}})$, the $D=4$ type II integral takes the form [^19] Recall the normalization $X_{I}\equiv X_{I}^{A}T^{A}$, $\mathrm{Tr}T^{A}T^{B}=\delta^{AB}$, where $T^{A}$ are $SU(N)$ or $U(N)$ generators. The measure $[dX]=\prod_{A=1}^{N^{2}-1}dX_{A}$ for $SU(N)$ or $[dX]=\prod_{A=1}^{N^{2}}dX_{A}$ for $U(N)$.

### v2

@@ 3 Classification of massive SYM matrix models
In this section, we provide the main idea that allows us to classify massive SYM matrix models. Here we will assume that we have one Majorana or Majorana-Weyl spinor, meaning $D=3,4,10$. In $D=6$, we have more than one Majorana spinor so we will treat this case separately (see appendix A.2). [^14] Note that the $D=2$ pure SYM matrix model does not admit any mass deformation. This is because one can neither write a fermionic mass term nor a Myers term in $D=2$. The constraints (24) and (25) then imply that all deformed quantities have to vanish.
Consider a massive deformation of $S_{0}$ (3) parameterized by a mass parameter $\mu$,
@@
while there are no (polynomial) combinations that can give lower-dimensional operators. [^15] Note that $\mathrm{Tr}X^{I}$ vanishes for traceless matrices. On the other hand, if one allows the matrices to have a trace, the term $\mathrm{Tr}X_{I}$ can always be removed by completing the square with the trace contribution coming from $S_{IJ}\mathrm{Tr}X_{I}X_{J}$. Inspired by Myers:1999ps, we refer to $S_{IJK}\mathrm{Tr}X^{I}X^{J}X^{K}$ as the Myers term.
The supersymmetries (4) may also need to be deformed in order to leave (15) invariant for all $\epsilon$. We consider the generic form
@@
Note that $\epsilon$ has mass dimension 1/2-1/2. Thus, $\delta_{B}^{(1)}X^{I},\delta_{B}^{(2)}X^{I},\ldots$ must be fermionic operators of dimension $1/2,-1/2,\ldots$. Since our fermions have dimension $3/2$, we cannot build such operators and thus $\delta_{B}^{(k)}=0$. On the other hand, $\delta_{F}^{(1)}\psi$ must be a bosonic operator of dimension 1. The only possibility is $\delta_{F}^{(1)}\psi\sim X_{I}$ while $\delta_{F}^{(k>1)}=0$. Putting everything together, the most general possibility at this point is
[…]
@@
[…]
@@
[…]
@@
This needs to be imposed for all spinors $\epsilon$, which are Majorana or Majorana-Weyl depending on the dimension. In order to proceed with solving these equations, one needs to specify the dimension $D$. The detailed analysis is carried out for each $D$ in appendix A. The resulting models are discussed in the next sections and summarized in table 1.
@@ 4 Ten-dimensional matrix model: polarized IKKT
The polarized IKKT model, originally introduced in Bonelli_2002 and recently examined in Hartnoll:2024csr; Komatsu:2024bop; Komatsu:2024ydh; Hartnoll:2025ecj, is a mass-deformed version of the IKKT model Ishibashi_1997 that retains invariance under sixteen (deformed) supersymmetries. We consider this model as the most interesting among all the matrix models, because of its richness as a holographic description of ten-dimensional geometries Komatsu:2024bop; Komatsu:2024ydh; Hartnoll:2024csr. In appendix A.1, we show that it is the unique [^16] Up to redefinition of $\mu$ and $SO(10)/(SO(3)\times SO(7))$ rotations. Moreover, we restrict to real bosonic deformations. mass deformation of IKKT that has sixteen supersymmetries and preserves $SU(N)$ gauge invariance.
@@
[…]
where $\epsilon$ is a Majorana-Weyl spinor (commuting) parameter, and $\delta_{\epsilon}$ is an anti-commuting (Grassmann-odd) operator. The supersymmetries together with the global $SO(3)\times SO(7)$ form either a gauged $F(4)$ or a gauged $F^{1}(4)$ superalgebra in Lorentzian signature. [^17] The anticommutator of two supersymmetries takes exactly the form expected from $F(4)$ as was shown in Komatsu:2024ydh. This model can be obtained from SYM on $S^{n}$ in the limit $n\to 0$ Blau:2000xg. [^18] In the notation of Blau:2000xg, polarized IKKT arises from the family B, with $(d=10,n=0)$. Note that as in IKKT, $g_{\mathrm{YM}}$ can be scaled out of the integral, by rescaling $X_{I}=\sqrt{g_{\mathrm{YM}}}\tilde{X}_{I}$ and $\psi_{\alpha}=g_{\mathrm{YM}}^{3/4}\tilde{\psi}_{\alpha}$. This leaves only $\Omega\equiv\mu/\sqrt{g_{\mathrm{YM}}}$ as a mass parameter.
@@
Let us now discuss observables, in dimensionless units where the only tunable parameter is $\Omega\equiv\mu/\sqrt{g_{\mathrm{YM}}}$. Firstly, the fixed $N$, $\Omega\to 0$ limit is accessible analytically (at leading order). In this case, one can integrate out the off-diagonal components. This allows us to compute a large set of gauge invariant observables in the strict massless limit as shown in (221). In particular, this analysis yields
@@
The large $\Omega$ behavior $\Omega\gg N^{1/4}$ was obtained in Hartnoll:2024csr yielding
@@ 5 Six-dimensional matrix models
The supersymmetry transformations take the form [^19] Recall that we use the convention where $\epsilon$ is a commuting parameter while $\delta_{\epsilon}$ is an anticommuting operator.
@@
In appendix A.2 we derive all the mass deformations of $S_{0}^{(D=6)}$ that carry eight supersymmetries as well as $SU(N)$ gauge invariance. We find two different models that we will denote by type I and type II. The type I model breaks $SO(6)\to SO(3)\times SO(3)$, and possesses a Myers term, so that the matrix integral has many different saddles as in polarized IKKT. The type II model has $SO(5)$ global symmetry and only has a trivial saddle.
@@ 5.1 Type I: $SO(6)\to SO(3)\times SO(3)$
Note that the $SU(2)$ symmetry $\psi_{a}\to U_{ab}\psi_{b}$ is preserved. In Lorentzian signature, the total global symmetry group would thus be $SO(2,1)\times SO(3)\times SU(2)$, and we have 8 supercharges. We thus expect the superalgebra to lie in the exceptional family $D(2,1,\alpha)$ VanProeyen:1999ni. This is indeed the case as we show in appendix A.2.1, where we find the specific value $\alpha=1/2$ which is not isomorphic to OSp(4|2)OSp(4|2). [^20] It is surprising that we did not find models with arbitrary values of $\alpha$, since $\alpha=1/2$ seems arbitrary. However, it is not hard to see by allowing the coefficients in (37) and (38) to take arbitrary values that supersymmetry is an algebraic constraint that leaves no free parameter except $\mu$. Thus, $\alpha=1/2$ is, surprisingly, the unique possibility. This model can be obtained from SYM on $S^{n}$ with 8 supersymmetries, extrapolating Blau:2000xg to $n=0$.  [^21] In the notation of Blau:2000xg, it arises from the family B, with $(d=6,n=0)$
@@
Finally, let us discuss observables in dimensionless coordinates $(g_{\mathrm{YM}}\to 1,\mu\to\Omega\equiv\mu/\sqrt{g_{\mathrm{YM}}})$. In the strict massless limit at fixed $N$, we can integrate out off-diagonals as shown in appendix C. This allows us to compute a large class of observables, including
@@ 5.2 Type II: $SO(6)\to SO(5)$
[…]
@@
Note that the $SU(2)$ symmetry $\psi_{a}\to U_{ab}\psi_{b}$ with $U\in SU(2)$ remains a symmetry only if $U^{\top}sU=s$ where $s$ is a two by two symmetric matrix with components $s_{12}=s_{21}=1$, $s_{11}=s_{22}=0$. This restricts it to a one parameter symmetry, infinitesimally equivalent to a $U(1)$. In Lorentzian signature, we thus have either $SO(5)\times U(1)$ or $SO(4,1)\times U(1)$ as the symmetry group, together with 8 supercharges. As we show in appendix A.2.2, the corresponding superalgebra is either gauged $OSp^{*}(2|4)$ or gauged $OSp^{*}(2|2,2)$ respectively. This model can be obtained from SYM on $S^{n}$ with 8 supersymmetries, extrapolating Blau:2000xg to n=0n=0.222222In the notation of Blau:2000xg, it arises from the family A, with $(d=6,n=0)$
Let us now discuss observables in dimensionless units $(g_{\mathrm{YM}}\to 1,\mu\to\Omega\equiv\mu/\sqrt{g_{\mathrm{YM}}})$. In the strict massless limit at fixed $N$, we can integrate out off-diagonals as shown in appendix C. This allows us to compute a large class of observables, including
@@ 6 Four-dimensional matrix models
where $\mathbf{X}_{I}$ are $(N^{2}-1)\times(N^{2}-1)$ matrices written in the adjoint representation. Namely, the original matrices are expanded as $X_{I}=X_{I}^{A}T^{A}$ where $T^{A}$ are the $SU(N)$ generators in the fundamental representation normalized by $\mathrm{Tr}T^{A}T^{B}=\delta^{AB}$. The structure constants are defined by $T_{A},T_{B}]=if_{ABC}T_{C}$. Finally, we define the $(N^{2}-1)\times(N^{2}-1)$ matrices $\mathbf{X}_{I}$, by the components $(\mathbf{X}_{I})_{AB}\equiv f_{ABC}X_{I}^{C}$. Thus, they are real and antisymmetric. In contrast to all other dimensions, it is important to note that this determinant is always positive semi-definite. The proof was first established in Ambjorn:2000bf. We review it in appendix B.
In appendix A.3, we derive all the mass deformations of the $D=4$ pure SYM matrix model that carry four supersymmetries as well as $SU(N)$ gauge invariance. We find two models that we denote by type I and type II. The type I model preserves $SO(4)$ symmetry, and is parameterized by a single mass parameter $\mu$. The type II model breaks $SO(4)\to SO(3)$, and it is parameterized by a continuous family of two mass parameters $\mu_{1},\mu_{2}$. It has a Myers term, thus the matrix integral has many saddles as in polarized IKKT. The Pfaffians get deformed to functions of $\mu$ (type I) or $\mu_{1}$ (type II). As long as $\mu,\mu_{1}\in\mathbb{R}$, the non-negativity of the Pfaffian is preserved.
@@ 6.1 Type I: $SO(4)\to SO(4)$
In Lorentzian signature, the symmetry group would be $SO(3,1)$, and there are 4 supercharges. In appendix A.3.1, we show that the resulting superalgebra is a (gauged) (complexified) direct sum of two simple Lie superalgebras $OSp(1|2)\oplus OSp(1|2)$. [^24] This is in the same spirit as $\mathfrak{so}(3,1)\cong\mathfrak{sl}(2,\mathbb{R})\oplus_{\mathbb{C}}\mathfrak{sl}(2,\mathbb{R})$. Surprisingly, this model cannot be obtained from SYM on $S^{n}$ Blau:2000xg, unlike the other matrix models.
@@
The normalization for the measure is explained in appendix E. The matrices $\mathbf{X}_{I}$ were defined below (55). As we show in appendix B, the Pfaffian is always positive semi-definite when $\Omega\in\mathbb{R}$. This implies that it can be written as the square root of a determinant. Moreover, this determinant can be simplified, yielding
@@
[…]
@@ 6.2 Type II: $SO(4)\to SO(3)$
[…]
@@
In addition to these symmetries, there are chiral rotations $\psi\to e^{i\alpha\Gamma_{*}}\psi$ that form a $U(1)$. In Lorentzian signature, the symmetry group is either $SO(3)\times U(1)$ or $SO(2,1)\times U(1)$, and there are 4 supercharges. In appendix A.3.2, we show that the corresponding superalgebra is either gauged $SU(2|1)$ or gauged $SU(1,1|1)$ respectively VanProeyen:1999ni. Two one-parameter subfamilies of this model can be obtained from SYM on $S^{n}$ with 4 supersymmetries, extrapolating Blau:2000xg to $n=0$. In the notation of Blau:2000xg, the one-parameter family $\mu_{1}=-\mu_{2}$ arises from family A with $(d=4,n=0)$, while the one-parameter family $\mu_{1}=3\mu_{2}$ arises from family B with $(d=4,n=0)$.
@@
In dimensionless units $(g_{\mathrm{YM}}\to 1,\mu_{i}\to\Omega_{i}\equiv\mu_{i}/\sqrt{g_{\mathrm{YM}}})$, the $D=4$ type II integral takes the form [^25] Recall the normalization $X_{I}\equiv X_{I}^{A}T^{A}$, $\mathrm{Tr}T^{A}T^{B}=\delta^{AB}$, where $T^{A}$ are $SU(N)$ or $U(N)$ generators. The measure $dX]=\prod_{A=1}^{N^{2}-1}dX_{A}$ for $SU(N)$ or $dX]=\prod_{A=1}^{N^{2}}dX_{A}$ for $U(N)$.